10. -10 points SerCP10 17.P049 If a battery is rated at 56.5 A h, how much total charge can it deliver before it goes dead?
Answer #1

56.5 Ah means it can deliver 56.5 A of current in time t = 1 h = 1*60*60 s = 3600 s

We know, I = Q/t

=> Q = It

= 56.5*3600

= 203,400 C

Kindly rate :)